A car is travelling at a 101 km/h with an acceleration of 5.1 m/s2. What was the speed 3.3 second earlier.

Round off the answer to 2 decimal places. Dont write any unit inside the answer box.
Answer #1

Using 1st kinematic equation:

V = U + a*t

U = Initial speed = ?

V = final speed = 101 km/hr = 101*(5/18) = 28.06 m/sec

a = acceleration of car = 5.1 m/sec^2

t = time taken = 3.3 sec

So,

U = V - a*t

U = 28.06 - 5.1*3.3

U = 11.23 m/sec = Speed of car 3.3 second earlier

Please Upvote.

If you need final answer in km/hr, then U = 11.23*18/5 = 40.43 km/hr